OpenStudy (anonymous):

You have just grduated from college and landed youyr first job, you purchase a house cost 250,000 and find a bank interest rate 6% that compounds monthly for 30 years. what is the monthly payment for this loan? what is the unpaid balance at the end of 5 years? what is the unpaid balance at the end of the 10th year? I think I should use the formula A=p(1+I)n but I am not sure please help.

OpenStudy (anonymous):

Yes yo do use that formula:\[C(t) = Co(1+r)^t\] Where, C(t) is the cost in the future Co is the initial cost r is the rate in decimals not % (6% = 0.06) t is the time period, in this case in years.

OpenStudy (anonymous):

I have A=250.000 (1+.06) 12 A=250.000 (1.06) 12 I don't know the next step

OpenStudy (anonymous):

mathmind can yo help me when you're done

OpenStudy (anonymous):

why did you put 12 as your t? @btm2011, yes i'll try

OpenStudy (anonymous):

I don't know should I put 30 cause thats the years

OpenStudy (anonymous):

hmm not quite. Here we are making a formula of the Cost over the period of time t so,\[C(t)=250,000(1+.06)^t\]\[C(t) = 250,000(1.06)^t\] this is the formula to find the cost at any time t (which is measure in years in this problem). If we want to know the unpaid balance at the end of 5 years, what do we do?

OpenStudy (anonymous):

ummm take 250,000 (1.06) 5 i divide then multiply

OpenStudy (anonymous):

what are you dividing ?

OpenStudy (anonymous):

250,000 with (1.06)

OpenStudy (anonymous):

that's a multiplication, isn't it?

OpenStudy (anonymous):

ok so 250,000 (1.06) I got 265 then I divide that by 5

OpenStudy (anonymous):

nope, do you know what a exponent is ? some number to the Nth power?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is that the 30 years

OpenStudy (anonymous):

oh heck no. that is not the formula at all for paying off a loan. that is the formula for interest.

OpenStudy (anonymous):

if i can figure out how to send an attachment i will send the correct formula.

OpenStudy (anonymous):

last formula is the one you use for loan payments.

OpenStudy (anonymous):

No, an exponent or power tells you how many times a number is multiplies by itself:\[2^2 = 2*2 = 4\]\[2^3 = 2*2*2 = 8\]\[2^4 = 2*2*2*2 = 16\] and so on.. so, what we have is:\[C(t)=250,000(1.06)^t\]in 5 years:\[C(5 years)=250,000(1.06)^{5years}\]\[C(5)= 250,000(1.06*1.06*1.06*1.06*1.06) = 334,556.39\]Hence, the unpaind balance would be: 334,556.39 - 25000 = 84556.4

OpenStudy (anonymous):

that's not it ? which one is it then?

OpenStudy (anonymous):

I got 8.833

OpenStudy (anonymous):

this is just not right. You are answering the following question: "if you invest $250,000 at 6% compounded monthly, how much will you have in 30 years" this question says something different. "What are your monthly payments?"

OpenStudy (anonymous):

@ Satellite73 yes and it then ask for the unpaid balance after 5 years and 10 years

OpenStudy (anonymous):

\[PMT=\frac{250000\times \frac{.06}{12}}{(1-(1+\frac{.06}{12})^{-12\times 30}}\]

OpenStudy (anonymous):

i got $1498.88 rounded. i can try again, but this sounds about right.

OpenStudy (anonymous):

sio take 250000 multiply by .06 then divide by 12

OpenStudy (anonymous):

as you can see the formula is a bit complicated so you have to be careful when entering it into a calculator. lots of room for mistake.

OpenStudy (anonymous):

mathmind, can you please help me

OpenStudy (anonymous):

do I divide .06 with 12 first then multiply the answer by 250000

OpenStudy (anonymous):

that is your numerator. order does not matter. multiply by 250000 by .06 then divide by 12 or divide .06 by 12 then multiply by 250000 get the same thing

OpenStudy (anonymous):

the denominator is much trickier. if you have a good calculator put it all in at once. be careful with parenthese

OpenStudy (anonymous):

I don't have one with me

OpenStudy (anonymous):

then you certainly cannot do this. you cannot compute these things by hand.

OpenStudy (anonymous):

I have a regular one not one that you are aking if I have

OpenStudy (anonymous):

what kind?

OpenStudy (anonymous):

just a regular made in china hybrid power one

OpenStudy (anonymous):

then you have to use lots of parentheses.

OpenStudy (anonymous):

ok I am tryn but just can't get this